OpenStudy (anonymous):

Re-write 2x - 3y is greater than or equal to 6 in slope-intercept format. Please help, I have no idea how to do this with inequalities :(

OpenStudy (texaschic101):

slope intercept form is y = mx + b 2x - 3y >= 6 subtract 2x from both sides -3y >= -2x + 6 now divide by -3, changing the sign y <= (-2/-3)x + 6/-3 y <= 2/3x - 2

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

thank you so much for helping. I think I get it now, I just wan't sure how to do it with the inequality sign.